Entries for Treasury Stock

On January 31, Wilderness Resorts Inc. reacquired 22,500 shares of its common stock at $31 per share. On April 20, Wilderness Resorts sold 12,800 of the reacquired shares at $40 per share. On October 4, Wilderness Resorts sold the remaining shares at $28 per share.

Journalize the transactions of January 31, April 20, and October 4. Refer to the Chart of Accounts for exact wording of account titles.

X

Chart of Accounts

CHART OF ACCOUNTS
Wilderness Resorts Inc.
General Ledger
ASSETS
110 Cash
120 Accounts Receivable
131 Notes Receivable
132 Interest Receivable
141 Merchandise Inventory
145 Office Supplies
151 Prepaid Insurance
181 Land
193 Equipment
194 Accumulated Depreciation-Equipment
LIABILITIES
210 Accounts Payable
221 Notes Payable
226 Interest Payable
231 Cash Dividends Payable
236 Stock Dividends Distributable
241 Salaries Payable
261 Mortgage Note Payable
EQUITY
311 Common Stock
312 Paid-In Capital in Excess of Par-Common Stock
315 Treasury Stock
321 Preferred Stock
322 Paid-In Capital in Excess of Par-Preferred Stock
331 Paid-In Capital from Sale of Treasury Stock
340 Retained Earnings
351 Cash Dividends
352 Stock Dividends
390 Income Summary
REVENUE
410 Sales
610 Interest Revenue
EXPENSES
510 Cost of Merchandise Sold
515 Credit Card Expense
520 Salaries Expense
531 Advertising Expense
532 Delivery Expense
533 Selling Expenses
534 Rent Expense
535 Insurance Expense
536 Office Supplies Expense
537 Organizational Expenses
562 Depreciation Expense-Equipment
590 Miscellaneous Expense
710 Interest Expense
Answer #1
Debit Credit
January 31 Treasury Stock 697500 =22500*31
        Cash 697500
April 20 Cash 512000 =12800*40
       Treasury Stock 396800 =12800*31
       Paid-In Capital from Sale of Treasury Stock 115200
October 4 Cash 271600 =(22500-12800)*28
Paid-In Capital from Sale of Treasury Stock 29100
       Treasury Stock 300700 =(22500-12800)*31